OpenStudy (anonymous):

on a map of uptown Charlotte , College street is perpendicular to Davidson Avenue. The equation y = -6x+8 representing College Street. What is the equation representing Davidson Avenue if it passes through the point ( 7 , -2 ) ?

OpenStudy (anonymous):

Can you tell me the slope of y=-6x+8?

OpenStudy (anonymous):

its -6x

OpenStudy (anonymous):

Just the number in front of x so -6 so what would the slope be of a line perpendicular to that?

OpenStudy (anonymous):

the oppsoite reciprocal : 1/6 (change the sign and flip the fraction)

OpenStudy (anonymous):

You then have to use slope-point form o f a line: y-y1 = m(x-x1) x1 = 7 y1=-2